The points (−5, 6) and (1, r) lie on a line with slope −3. Find the missing coordinate r.
-Please help me..I have no idea what to do.

Answer by lwsshak3(11628)

The points (−5, 6) and (1, r) lie on a line with slope −3. Find the missing coordinate r.
***
Equation for a line: y=mx+b, m=slope, b=y-intercept
Using coordinates of given points (-5,6) and (1,r) on the line:
slope,m=∆y/∆x=(r-6)/(1-(-5))=(r-6)/(1+5)
(r-6)/(6)=-3
r-6=-18
r=-12